# Discrete Neural Algorithmic Reasoning

### v2

@@ 4 Experiments
Also, we are interested if discrete neural reasoners can be learned without hints and how they tend to utilize the given amount of node and edge states to solve the problem. We discuss no-hint experiments separately in Section 6.
@@ 4.1 Datasets
We perform our experiments on the problems from the recently proposed SALSA-CLRS benchmark (Minder et al., 2023), namely BFS, DFS, Prim, Dijkstra, Maximum Independent Set (MIS), and Eccentricity. We believe that the proposed method is not limited by the covered problems, but we leave the implementation of the required data flows (e.g., edge-based reasoning (Ibarz et al., 2022), graph-level hints, interactions between different scalars) for future work.
The train dataset of SALSA-CLRS consists of random graphs with at most 16 nodes sampled from the Erdös-Renyi (ER) distribution with parameter $p$ chosen to be as low as possible while graphs remain connected with high probability. The test set consists of sparse graphs of sizes from 16 to 1600 nodes.
We slightly modify the hints from the benchmark without conceptual changes (e.g., we have modified the hints for the DFS problem to remove graph-level hints). Discrete states are fully described by the non-scalar hints and scalars are exactly the hints of the $scalar$ type (we refer to Veličković et al. (2022) for the details on hint design).
@@ 4.2 Baselines and evaluation
We compare the performance of our proposed discrete model with two baseline sparse models, GIN (Xu et al., 2019) and Pointer Graph Network (Veličković et al., 2020a). We report both node-level and graph-level metrics for the baselines and our model. Also, we compare our model with Triplet-GMPNN (Ibarz et al., 2022) and two recent approaches, namely Hint-ReLIC (Bevilacqua et al., 2023) and G-ForgetNet (Bohde et al., 2024), which demonstrate state-of-the-art performance of hint-based neural algorithmic reasoning. However, as these methods are evaluated on the CLRS-30 benchmark and their code is not yet publicly available, we can only compare them on the corresponding tasks (BFS, DFS, Dijkstra, Prim) and CLRS-30 test data, namely ER graphs with $p=0.5$ of size 64. Note that this test data is more dense than that of SALSA-CLRS, meaning shorter roll-outs for given tasks. Also, only node-level metrics have been reported for these methods.
@@ 4.3 Model details
For our experiments, we use the model described in Section 3. We use one attention head for each scalar value. The number of processor steps is defined externally as the length of the ground truth algorithm trajectory, which is consistent with the prior work. We use one architecture (except the task-depended encoders/decoders), including the $\mathrm{ScalarUpdate}$ module for all the problems.
We recall that neural reasoners can operate either on the base graph (which is defined by the problem) or on more dense graphs with the original graph encoded into edge features. SALSA-CLRS proposes to enhance the size-generalization abilities of neural reasoners with increased test sizes (up to 100-fold train-to-test scaling), so we use the base graph for message-passing, similar to the SALSA-CLRS baselines. We also add a virtual node that communicates with all nodes of the graph.
@@ 4.4 Training details and hyperparameters
We train each model using the Adam optimizer, learning rate $\eta=0.001$, using teacher forcing, batch size of 32 graphs with 1000 optimization steps, and evaluate the resulting model. We anneal softmax temperatures in the discrete bottlenecks (attention weights, $\mathrm{ScalarUpdate}$ operations) geometrically from 3.0 to 0.01, decreasing the temperature at each training step. We report all hyperparameters in the source code. [^1] https://github.com/yandex-research/dnar
[…]
@@
[…]
Ecc. 16 25.3 100. 100.
80 23.8 100. 100.
160 26.1 100. 100.
800 17.1 100. 100.
1600 16.0 83.0 100.
During training, we minimize the standard hints and output losses: scalar hints are optimized with MSE loss, and other types of hints are optimized with cross-entropy and categorized cross-entropy losses (Veličković et al., 2022; Ibarz et al., 2022). Note that we do not supervise any additional details in model behavior, e.g., selecting the most important neighbor in the attention block, the exact operations with scalars, etc.
For multitask experiments, we follow the setup proposed by Ibarz et al. (2022) and train a single processor with task-dependent encoders/decoders. We make 10000 optimization steps on the accumulated gradients across each task and keep all hyperparameters the same as in the single task.
@@ 4.5 Results
To summarize, we have two key components of our contribution: separating discrete and continuous data flows and discrete bottlenecks. Note that using discrete bottlenecks without the separation between flows is not applicable, as would lead to the loss of the continuous information, as we mention in Section 3. On the other hand, we ablate the discretization block to verify its contribution to the performance of the proposed model. Our experiments show that the proposed processor without the discretization performs non-above the baselines level on all the tasks.
@@
Task Triplet-GMPNN Hint-ReLIC G-ForgetNet DNAR (ours)
BFS 99.73 $\pm$ 0.0 99.00 $\pm$ 0.2 99.96 $\pm$ 0.0 100. $\pm$ 0.0
DFS 47.79 $\pm$ 4.2   100. $\pm$ 0.0 74.31 $\pm$ 5.0 100. $\pm$ 0.0
Dijkstra 96.05 $\pm$ 0.6 97.74 $\pm$ 0.5 99.14 $\pm$ 0.1 100. $\pm$ 0.0
MST-Prim 86.39 $\pm$ 1.3 87.97 $\pm$ 2.9 95.19 $\pm$ 0.3 100. $\pm$ 0.0
@@ 5 Interpetability and testing
In addition to the empirical evaluation of the trained models on diverse test data, the discrete and size-independent design of the proposed models allows us to interpret and test them manually. The main idea is to show that the model will perform the exact discrete state transitions (including discrete operations with scalars) as the original algorithm.
First, we note that due to the hard attention, each node receives exactly one message. Also, the message depends only on the discrete states of the corresponding nodes and edges. Thereby, as each node and edge states after a single processor step depend only on the current states and received message, all the possible options can be directly enumerated and tested if all states change to the correct ones.
The only remaining part to fully interpret the whole model is the attention block. We note that our implementation of the $select\_best$ selector (Section 3.3) does not necessarily produce the top-1 choice over the ordered pairs of ‘states priority’ and scalars $s_{ij}$ as it simply augments key vectors with indicators if the given edge has the best scalar among others. For example, it may happen that for some state, the maximum attention weight is achieved for an edge without the indicator. However, given the finite number of discrete states, we can manually check if the mentioned “best” indicator increases the attention weight between every pair of states. Combined with hard attention, this would imply that the attention block attends depending on the predefined states and uses scalar priorities only to break ties.
Given that, we can unit-test all possible state transitions and attention blocks. With full coverage of such tests, we can guarantee the correct execution of the desired algorithm for any test data. We tested our trained models from Section 4 manually verifying state transitions. As a result, we confirm that the attention block indeed operates as $select\_best$ selector, as the model actually uses these indicators to increase the attention weights. Thus, we can guarantee that for any graph size, the model will mirror the desired algorithm, which is correct for any test size.
@@ 6 Towards no-hint discrete reasoners
Training deep discretized models is known to be challenging, therefore we focus only on the BFS algorithm, as it is well-aligned with the message-passing framework, has short roll-outs, and can be solved with small states count (note that for no-hint models node/edge states count is a hyperparameter due to the absence of the ground truth states trajectory).
@@
We perform hyperparameters search over training graph sizes, discrete states count, softmax temperature annealing schedules, and training steps. The trained models never achieved the perfect validation scores. Then, we selected the best-performing models and tried to reverse-engineer how they utilize the given states. In short, we found that the model operating on $k$ node states tends to use them to assign the layers from the starting node, except for the nodes on a distance greater than $k$, where state updates do not mirror the expected BFS flow. Also, the major part of correctly predicted pointers is concentrated on the first $k$ layers. We found this behavior well-aligned to the BFS algorithm and indicating the possibility of achieving the perfect validation score with enough states count. However, this algorithm does not generalize since it fails at distances larger than those encountered during training.
On the other hand, one can demonstrate that for a small enough state count (for BFS, it is two node states and two edge states) and diverse enough validation data, the perfect validation performance implies that the learned solution will generalize to any graph size.
@@ 7 Limitations and future work
First, several proposed design choices strictly reduce the expressive power of the model. For example, due to the hard attention, the proposed model is unable to compute the mean value from all the neighbors in a single message-passing step, which is trivial for attention-based models (note that this can be computed in several message-passing steps). Thus, the model in its current form is unable to express transitions between hints for some algorithms from the CLRS-30 in a single processor step. However, we believe that the expressivity of the proposed model can be enhanced with additional architectural modifications (e.g., edge-based reasoning (Ibarz et al., 2022), global states, interactions between different scalars) that can be combined with the proposed discretization ideas.
Second, while we report the perfect scores for the covered tasks, we cannot guarantee that the training will converge to the correct model for any initialization/training data distributions. However, we empirically found the proposed method to be quite robust to various architecture/training hyperparameter choices.
@@
Our method is based on the particular architectural choice and actively utilizes the attention mechanism. However, the graph deep learning field is rich in various architectures exploiting different inductive biases and computation flows. The proposed separation between discrete states and continuous inputs may apply to other models, however, any particular construction can require additional efforts.
Also, we provide only one example of the $\mathrm{ScalarUpdate}$ block. We believe that utilizing a general architecture (e.g., some form of discrete Neural Turing Machine (Graves et al., 2014; Gulcehre et al., 2016)) capable of executing a wider range of manipulation is of interest for future work.
With the development of neural reasoners and their ability to execute classic algorithms on abstract data, it is becoming more important to investigate how such models can be applicable in real-world scenarios according to the Neural Algorithmic Reasoning blueprint (Veličković & Blundell, 2021) and transfer their knowledge to high-dimensional noisy data with intrinsic algorithmic dependencies. While there are several examples of NAR-based models tackling real-world problems (Beurer-Kellner et al., 2022; Numeroso et al., 2023), there are no established benchmarks for extensive evaluation and comparison of different approaches.
Lastly, we leave for future work a deeper investigation of learning interpretable neural reasoners without hints, which we consider essential from both theoretical perspective and practical applications, e.g., combinatorial optimization.

### v3

@@ 4 Experiments
Additionally, we are interested in whether discrete neural reasoners can be learned without hints and how they tend to utilize the given number of node and edge states to solve the problem. We discuss no-hint experiments separately in Section 6.
Finally, we investigate whether graph sizes in the train set can be reduced without compromising generalization performance; see Appendix B for details.
@@ 4.1 Datasets
We perform our experiments on the problems from the recently proposed SALSA-CLRS benchmark (Minder et al., 2023), namely BFS, DFS, Prim, Dijkstra, Maximum Independent Set (MIS), and Eccentricity. We believe that the proposed method is not limited to the covered problems, but we leave the implementation of the required data flows (e.g., edge-based reasoning (Ibarz et al., 2022), graph-level hints, interactions between different scalars) for future work.
The train dataset of SALSA-CLRS consists of random graphs with at most 16 nodes sampled from the Erdös-Rényi (ER) distribution with parameter $p$ chosen to be as low as possible such that graphs remain connected with high probability. The test set consists of sparse graphs of sizes from 16 to 1600 nodes.
We slightly modify the hints from the benchmark without conceptual changes (e.g., we have modified the hints for the DFS problem to remove graph-level hints and node discovery/finish times). Our discrete states are defined by the benchmark’s non-scalar hints, while our scalars are exactly the hints of the scalarscalar type (see Veličković et al. (2022) for details on hint design).
@@ 4.2 Baselines and Evaluation
We compare the performance of our proposed discrete model with two baseline sparse models, GIN (Xu et al., 2019) and Pointer Graph Network (Veličković et al., 2020a). We report both node-level and graph-level metrics for the baselines and our model. Graph-level metrics measure the proportion of test graphs for which all node-level outputs are predicted correctly. For example, for the BFS problem, the output of the algorithm is a tree, where each node points to its parent in the BFS exploration tree (and the starting node points to itself). Pointer is a specific hint/output type which forces each node to point to exactly one node.
Additionally, we compare our model with Triplet-GMPNN (Ibarz et al., 2022) and two recent approaches, namely Hint-ReLIC (Bevilacqua et al., 2023) and G-ForgetNet (Bohde et al., 2024), which demonstrate state-of-the-art performance in hint-based neural algorithmic reasoning. However, as these methods are evaluated on the CLRS-30 benchmark and their code is not yet publicly available, we only compare them on the corresponding tasks (BFS, DFS, Dijkstra, Prim) and the CLRS-30 test data, namely ER graphs with $p=0.5$ of size 64. Note that this test data is denser than that of SALSA-CLRS, meaning shorter roll-outs for the given tasks. Furthermore, only node-level metrics have been reported for these methods.
@@ 4.3 Model Details
For our experiments, we use the model described in Section 3. We use one attention head for each scalar value. The number of processor steps is set to the length of the ground-truth algorithm trajectory, which is consistent with prior work. We use one architecture (except the task-dependent encoders/decoders), including the $\mathrm{ScalarUpdate}$ module, for all the problems.
We recall that neural reasoners can operate either on the original graph (which is defined by the problem) or on denser graphs with the original graph encoded into edge features. SALSA-CLRS proposes to enhance the size-generalization abilities of neural reasoners with increased test sizes (up to 100-fold train-to-test scaling), so we use the original graph for message-passing, similar to the SALSA-CLRS baselines. We also add a virtual node that communicates with all nodes of the graph.
@@ 4.4 Training Details and Hyperparameters
We train each model for 1000 optimization steps using the Adam optimizer with a learning rate of $\eta=0.001$ and a batch size of 32 graphs; teacher forcing is applied throughout training. For the discrete bottlenecks (attention weights and $\mathrm{ScalarUpdate}$ operations), we anneal the softmax temperature geometrically from 3.0 to 0.01, decreasing it at each training step. A complete list of all hyperparameters is provided in our source code. [^2] https://github.com/yandex-research/dnar
During training, we minimize the standard hint and output losses: scalar hints are optimized with the MSE loss, and the other types of hints are optimized with the cross-entropy and categorized cross-entropy losses (Veličković et al., 2022; Ibarz et al., 2022). Notably, we do not supervise any additional details in model behavior, such as selecting the most important neighbor in the attention block or the exact operations with scalars.
Table 1: Node/graph-level test scores for the proposed discrete reasoner and the baselines on SALSA-CLRS test data. Scores are averaged across 5 different seeds; standard deviation is omitted. For the eccentricity problem, only the graph-level metric is applicable.
@@
BFS 16 98.8 / 92.5 100. / 100. 100. / 100.
80 95.3 / 59.4 99.8 / 88.1 100. / 100.
160 95.1 / 37.8 99.6 / 66.3 100. / 100.
800 86.9 / 0.9 98.7 / 0.2 100. / 100.
1600 86.5 / 0.0 98.5 / 0.0 100. / 100.
DFS 16 41.5 / 0.0 82.0 / 19.9 100. / 100.
80 30.4 / 0.0 38.4 / 0.0 100. / 100.
160 20.0 / 0.0 26.9 / 0.0 100. / 100.
800 19.5 / 0.0 24.9 / 0.0 100. / 100.
1600 17.8 / 0.0 23.1 / 0.0 100. / 100.
SP 16 95.2 / 49.8 99.3 / 89.5 100. / 100.
80 62.4 / 0.0 94.2 / 3.3 100. / 100.
160 53.3 / 0.0 92.0 / 0.0 100. / 100.
800 40.4 / 0.0 87.1 / 0.0 100. / 100.
1600 36.9 / 0.0 84.5 / 0.0 100. / 100.
Prim 16 89.6 / 29.7 96.4 / 69.9 100. / 100.
80 51.6 / 0.0 79.7 / 0.0 100. / 100.
160 49.5 / 0.0 75.6 / 0.0 100. / 100.
800 45.0 / 0.0 69.5 / 0.0 100. / 100.
1600 43.2 / 0.0 66.8 / 0.0 100. / 100.
MIS 16 79.9 / 3.3 99.8 / 98.6 100. / 100.
80 79.9 / 20.0 99.4 / 88.9 100. / 100.
160 78.2 / 0.0 99.4 / 76.2 100. / 100.
800 83.4 / 0.0 98.8 / 18.0 100. / 100.
1600 79.2 / 0.0 98.9 / 5.2 100. / 100.
Ecc. 16 NA / 25.3 NA / 100. NA / 100.
80 NA / 23.8 NA / 100. NA / 100.
160 NA / 26.1 NA / 100. NA / 100.
800 NA / 17.1 NA / 100. NA / 100.
1600 NA / 16.0 NA / 83.0 NA / 100.
For multitask experiments, we adopt the setup from Ibarz et al. (2022) and train a single processor with task-dependent encoders/decoders to imitate all covered algorithms simultaneously. We make 10000 optimization steps on the accumulated gradients across each task and keep all hyperparameters the same as in the single task.
@@ 4.5 Results
Recall that we have three key components of our contribution: feature discretization, hard attention, and separating discrete and continuous data flows. To evaluate the importance of each component for generalization capabilities of the proposed models, we conduct an ablation study; the details can be found in Appendix A. In short, we demonstrate that removing each of these components yields the model without provable guarantees of perfect generalization. However, these components differ in terms of the impact on the performance. In particular, using regular attention instead of hard attention yields perfect test scores for given datasets, but it is possible to construct adversarial examples with large neighborhood sizes where performance drops. On the other hand, removing discretization from the scalar updater significantly affects the performance even on the small test graphs.
@@
Task Hint-ReLIC G-ForgetNet DNAR (ours)
BFS 99.00 $\pm$ 0.2 99.96 $\pm$ 0.0 100. $\pm$ 0.0
DFS   100. $\pm$ 0.0 74.31 $\pm$ 5.0 100. $\pm$ 0.0
SP 97.74 $\pm$ 0.5 99.14 $\pm$ 0.1 100. $\pm$ 0.0
Prim 87.97 $\pm$ 2.9 95.19 $\pm$ 0.3 100. $\pm$ 0.0
@@ 5 Interpetability and Testing
In addition to the empirical evaluation of the trained models on diverse test data, the discrete and size-independent design of the proposed models allows us to interpret and test them manually. Namely, we can validate that the model will perform the exact discrete state transitions (including discrete operations with scalars) as the original algorithm.
First, due to the hard attention, each node receives exactly one message. Also, the message depends only on the discrete states of the corresponding nodes and edges. Thus, node and edge states after a single processor step depend only on the current states and the received message. Therefore, all the possible options can be directly enumerated to test if all states change to the correct ones.
The final component to fully interpret the whole model is the attention block. We note that our implementation simply augments the key vectors with the indicators of whether the given edge has the best scalar among others. For example, it may happen that for some state, the maximum attention weight is achieved for an edge without the indicator. However, given the finite number of discrete states, we can manually check the unnormalized attention scores between each pair of states both with this indicator and without it. Combined with hard attention, this would imply that the attention block attends to nodes based on their predefined states and uses scalar priorities only to break ties. A particular example with a more detailed explanation for the BFS problem can be found in Appendix D.
Given that, we can unit-test all possible state transitions and attention blocks. With full coverage of such tests, we can guarantee the correct execution of the desired algorithm for any test data. We tested our trained models from Section 4 by manually verifying state transitions. As a result, we confirm that the attention block indeed operates as select_bestselect\_best selector, as the model actually uses these indicators to increase the attention weights. Thus, we can guarantee that for any graph size, the model will mirror the desired algorithm, which is correct for any test size.
@@ 6 Towards No-Hint Discrete Reasoners
Training deep discretized models is known to be challenging (Mahdavi et al., 2023). We observe that without hyperparameter search, discrete models are only slightly improved over the untrained models. Therefore, we focus only on the BFS algorithm, as it is well-aligned with the message-passing framework, has short roll-outs, and can be solved with a small number of states (note that for no-hint models node/edge state count is a hyperparameter due to the absence of the ground-truth states trajectory).
@@
We perform hyperparameter search over graph sizes in the train set (using ER graphs with $p=0.5$ and $n\in[4,16]$), discrete node state count (from 2 to 6 states), softmax temperature annealing schedules $([3,0.01],[3,0.1],[3,1])$. For each hyperparameter choice, we train 5 models with different seeds. We validate the resulting models on the graphs of size 16. The best resulting model is obtained with the graph size $n=5$ and 4 node states. The trained models never achieved perfect validation scores (see Table 3 for the results).
Table 3: BFS node/graph-level scores of the best_no_hint_model for different graph sizes.
5 16 64
best_no_hint_model 99 / 86 94 / 34 79 / 0
Then, we select the best-performing models and try to analyze the errors of the resulting models and reverse-engineer how they utilize the given states. First, we look at the node states after the last step of the processor and note that the states correspond to the distances from the starting node. More formally, we note that the model with four states uses the first state for the starting node, the second state for its neighbors, the third state for nodes at distance two from the starting node, and the last state for all other nodes. Such states-based classification of distance has accuracy >> 98% when tested on 1000 random graphs with 16 nodes. We observe that for the nodes that belong to the first four distance layers from the starting node, the pointers are predicted with 100% accuracy and these pointers are computed layer-by-layer as in the ground-truth algorithm (we refer to Appendix E for illustrations). The model’s errors occur at distances $\geq 4$ from the starting node (we did not reverse-engineer the specific logic of computations on larger distances).
We found this behavior well-aligned with the BFS algorithm, which indicates the possibility of achieving the perfect validation score with enough number of states. However, this algorithm does not generalize since it fails at distances larger than those encountered during training.
On the other hand, one can demonstrate that for a small enough state count (for BFS, it is two node states and two edge states) and sufficiently diverse validation data, the perfect validation performance implies that the learned solution will generalize to any graph size.
@@
Finally, we note that for sequential problems, such as DFS, obtaining a good no-hint model can be even more challenging and may require additional effort. One possible way to overcome this limitation is to implement a curriculum learning setup, which we leave for future work.
@@ 7 Limitations and Future Work
First, several proposed design choices strictly reduce the expressive power of the model. For example, due to the hard attention, the proposed model is unable to compute the mean value from all the neighbors in a single message-passing step, which is trivial for attention-based models (note that this can be computed in several message-passing steps). Thus, the model in its current form is unable to express transitions between hints for some algorithms in the CLRS-30 benchmark in a single processor step.
However, we believe that the expressivity of the proposed model can be enhanced with additional architectural modifications (e.g., edge-based reasoning (Ibarz et al., 2022), global states, interactions between different scalars) that can be combined with the proposed discretization ideas. Additionally, there are several potential ways to balance between expressive power of the model and strong generalization guarantees, e.g., removing hard attention or removing feature discretization (while keeping the discretization inside the $\mathrm{ScalarUpdate}$ module).
Second, while we report perfect scores for the covered tasks, we cannot guarantee that the training will converge to the correct model for any initialization/training data distributions. However, we empirically found the proposed method to be quite robust to various architecture/training hyperparameter choices.
@@
Our method is based on the particular architectural choice and actively utilizes the attention mechanism. However, the graph deep learning field is rich in various architectures exploiting different inductive biases and computation flows. The proposed separation between discrete states and continuous inputs may apply to other models. However, any particular construction may require additional effort.
Also, we provide only one example of the $\mathrm{ScalarUpdate}$ block. We believe that utilizing a general architecture (e.g., some form of discrete Neural Turing Machine (Graves et al., 2014; Gulcehre et al., 2016)) capable of executing a wider range of manipulations is of interest for future work.
With the development of neural reasoners and their ability to execute classical algorithms on abstract data, it is becoming more important to investigate how such models can be applicable in real-world scenarios according to the Neural Algorithmic Reasoning blueprint (Veličković & Blundell, 2021) and transfer their knowledge to high-dimensional noisy data with intrinsic algorithmic dependencies. While there are several examples of NAR-based models tackling real-world problems (Beurer-Kellner et al., 2022; Numeroso et al., 2023), there are no established benchmarks for extensive evaluation and comparison of different approaches.
Lastly, we leave for future work a deeper investigation of learning interpretable neural reasoners without hints, which we consider essential both from a theoretical perspective and for practical applications, e.g., combinatorial optimization.